13. The energy state with n=1 has the energy of (a) +13.6 eV (b) -1.51eV (c)- (d) N
Answer #1

Energy of nth state is given by : En a - 13.6 evi n2 for(n=1) - E,= -13.6 (۱) - ...- E =-13.6ev .